OpenStudy (anonymous):

Solve the equation Quantity 7 x minus 17 all over 8 = 4 for x. Please show your work and solution in the space provided.

OpenStudy (anonymous):

\[\frac{7 x - 17}{8} = 4 x. \rightarrow 7 x - 17 =32x \rightarrow 7 x - 32x=-17 \] \[-25x=-17 \rightarrow x= 17/25\]

OpenStudy (anonymous):

Thank you so much!